Al sold a total of 300 Pizzas and Gyros. Pizzas were sold at $8 each and Gyros for $6. Total sales came to $2,000. How many of each was sold?
Answer by checkley77(12844)

P+G=300 OR P=300-G
8P+6G=2,000
8(300-G)+6G=2,000
2,400-8G+6G=2,000
-2G=2,000-2,400
-2G=-400
G=-400/-2
G=200 GYROS WERE SOLD.
P=300-200
P=100 PIZZA WERE SOLD.
PROOF:
8*100+6*200=2,000
800+1,200=2,000
2,000=2,000